OpenStudy (anonymous):

Multiply and answer in scientific notation: (3.8 x 10^8)(6.9 x 10^5) A) 2.622 x 10^13 B) 2.622 x 10^14 C) 26.22 x 10^13 D) 26.22 x 10^14

OpenStudy (anonymous):

@mathsciencehistory

OpenStudy (anonymous):

Ill be right back!

OpenStudy (jhannybean):

\[\large (3.8\cdot 6.9) \times 10^{8+5}\]

OpenStudy (anonymous):

im about to have to go sorry

OpenStudy (anonymous):

I would say the answer is A or D

OpenStudy (anonymous):

A

OpenStudy (jhannybean):

Solve for my expression and you will find your answer.

OpenStudy (mathstudent55):

The correct answer hasn't been mentioned yet.

OpenStudy (anonymous):

c

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

Im back!

OpenStudy (anonymous):

So C

OpenStudy (jhannybean):

Good job.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

me??

OpenStudy (anonymous):

Im sorry i cant give out medals but i fanned you

OpenStudy (anonymous):

i know

OpenStudy (mathstudent55):

\( (3.8 \times 10^8)(6.9 \times 10^5)\) \(= (3.8 \times 6.9) \times (10^8 \times 10^5)\) \(= 26.22 \times 10^{8 + 5} \) \(= 26.22 \times 10^{13}\) The number above is the correct product, but it is not in scientific notation. \(= 2.622 \times 10^{14} \) Now it is correct and in scientific notation.